OpenStudy (anonymous):

how can i solve 5(x+1) + 2(x+3) = 0?

OpenStudy (anonymous):

Multiply the terms inside the brackets, then open then and equate to 0.

OpenStudy (anonymous):

so 5x + 5 + 2x + 6 = 0 7x + 11 = 0 7x = -11 after this?

OpenStudy (anonymous):

Find x by taking 7 in the denominator of the other number i.e. -11.

OpenStudy (anonymous):

x = -11/7 is this correct

OpenStudy (anonymous):

great, well done.

OpenStudy (anonymous):

thnx alot!